OpenStudy (anonymous):

What is the value of the expression? 19.2 + (–7 + 3.8) A. 16 B. 24 C. 29 D. 30

OpenStudy (zale101):

the order of operation states that u first need to solve the parenthesis

OpenStudy (anonymous):

A 16 You start with the Parenthesis -7+3.8 gives you -3.2 Then take 19.2 and add -3.2 which is 16

OpenStudy (zale101):

PEMDAS Please Excuse MY Dear Aunt Sally Parenthesis, exponent, multiplication, division, addition, subtraction

OpenStudy (anonymous):

Thank you :) guys

OpenStudy (anonymous):

Help me with one more?

OpenStudy (anonymous):

Sure. Go for it.

OpenStudy (anonymous):

What value of q makes the statement true? –(q + 14) = –3 + (–14) A. –17 B. –3 C. 3 D. 17

OpenStudy (anonymous):

So first off you need to start with parenthesis again. Since you cant do anything with (q+14) you wanna distribute the - which then turn it into -Q-14=-3+(-14) This equation turns into -q-14=-3-14 because if you add a negitive you subtract. Then you want to isolate the variable -q. You don't want the variable to be negative so add Q to the left and right side of the equation and then do the same with -3 and -14 add them to both sides. You will end with 3=q

OpenStudy (anonymous):